A rock is thrown upward from level ground in such a way that the maximum height of its flight is equal to its horizontal range R. At what angle u is the rock thrown?
1
Expert's answer
2021-02-19T10:32:32-0500

The vertical and horizontal components are, respectively, "v\sin \alpha, v\cos\alpha". The time to the highest point (where the vertical component vanishes) is "T=\frac{v\sin \alpha}{g}", and by symmetry the total time until the rock falls is "t=2T = \frac{2v\sin\alpha}{g}" . The maximum height attained can be expressed as "h=\frac{v^2\sin^2\alpha}{2g}" (for a uniformly accelerated movement we have "d= \frac{v_{final}^2-v_{initial}^2}{2a}", where "v_{final},v_{initial}, a" are taken in the direction of movement). The horizontal range can be expressed as "R=v\cos \alpha \cdot t=\frac{2v^2\sin\alpha \cos\alpha}{g}", as in the horizontal direction the movement is uniform (we neglect the air resistance etc). Therefore, "R=h" is achieved when "2\sin\alpha \cos \alpha=\frac{\sin^2\alpha}{2}". This equation has two solutions (we take "0\leq \alpha \leq \pi/2") : "\alpha=0" (the obvious one, if the rock is thrown parallel to the ground from the level of the ground, it does not travel at all) and "\tan \alpha = 4", "\alpha=\arctan 4 \approx 76^\circ".